# Work extraction with feedback control using limited resources

### v1

@@ Abstract
Many physical, biological, and even social systems are faced with the problem of how to efficiently harvest free energy from any of many possible environmental states, yet only have a limited number of harvesting strategies to choose among. We formalize this general situation within the framework of work extraction from nonequilibirum systems using feedback control. Specifically, we analyze the combined and separate effects of noisy measurements, memory limitations, and the access to only a limited set of possible control operations, quantifying the amount of extractable work under all such constraints. Such analysis allows us to provide a general recipe to construct repertoires of allowed control operations that minimize the expected losses during the feedback protocol, i.e., that minimize expected entropy production. Our results highlight the importance and benefits of feedback control over uninformed (random) actions beyond the information gain under control restrictions and are applied to several situations including uncertain initial distributions.
@@ I Introduction
The acquisition and processing of information has been recognized as a crucial element that can aid functionality and gain thermodynamic performance for natural organisms and small synthetic devices that harvest free energy in fluctuating environments. This can facilitate a wide range of biological functions and the performance of thermodynamic tasks such as power generation or refrigeration.
@@
Since Maxwell’s demon thought experiment in the 19th century and the first information engine using feedback control operations proposed by Leo Szilard Szilard (1929), a long tradition of scientists have been fascinated by the interplay between the probabilistic nature of the second law and the possibility to use information to improve thermodynamic operation Leff and Rex (2002). The development of stochastic thermodynamics in the last decades has provided the possibility of a precise formulation and analysis of feedback control scenarios Parrondo et al. (2015) where the costs and benefits of using information can be assessed in a variety of situations, ranging from discrete measurements Touchette and Lloyd (2000); Sagawa and Ueda (2009, 2010); Horowitz and Vaikuntanathan (2010), to continuous monitoring Sagawa and Ueda (2012); Deffner and Jarzynski (2013); Potts and Samuelsson (2018); Ribezzi-Crivellari and Ritort (2019); Debiossac et al. (2020), and from classical to quantum systems Sagawa and Ueda (2008); Funo et al. (2013); Francica et al. (2017); Manzano et al. (2018); Annby-Andersson et al. (2022). Experiments in a broad range of platforms from colloidal particles Toyabe et al. (2010) and nanoelectronic devices Koski et al. (2014a, b) to DNA pulling experiments Rico-Pasto et al. (2021) have tested the role of information in work extraction.
In this work, we focus on limitations in work extraction by feedback protocols under noisy observation that arise from a limited number of actions (i.e. memory size limitations) and from the quality of the actions that the feedback controller can perform. Such limitations arise in many experimental scenarios where the set of effective actions the controller can implement to extract work is reduced to a few operations. For example, switching on and off predefined potentials Toyabe et al. (2010), suddenly raising the energy of a discrete level Koski et al. (2014a) or swapping the polarity of photons Vidrighin et al. (2016). Under such limitations, the possibility of feedback control gains new value beyond the thermodynamic value of the acquired information: it enables work extraction in situations in which it would be otherwise impossible. However, these action limitations can also spoil or drastically reduce the efficiency of work extraction itself, as we show. As a result, optimizing the set of actions subject to memory limitations, input distributions, and, noise in measurements is crucial in several applications. These include cases where the initial distribution of the system is not known and might be guessed from noisy observation.
@@
To analyze the consequences of these limitations, we build on previous work demonstrating that optimal feedback control operation which maximizes work extraction can be achieved with reversible feedback protocols Horowitz and Parrondo (2011). This can be done by combining a quench that depends on the initial system distribution and the measurement results with a quasi-static isothermal processes Hasegawa et al. (2010); Esposito and den Broeck (2011); Parrondo et al. (2015). However, in real-world scenarios, limitations often arise due to both imperfections in the measurements and because of severe constraints in the control operations that prevent the implementation of optimal protocols Johan du Buisson and Bechhoefer (2024). For example, if the relaxation timescales of the system are not fast enough with respect to the driving velocity, the quasi-static regime is lost. In that case, optimal transport techniques are required to maximize the extractable work that usually involve numerical techniques Aurell et al. (2011). In some cases, however, the problem can be directly solved in a linear response regime if the dynamics is sufficiently slow (i.e. close to but out of the quasi-static regime) Sivak and Crooks (2012); Zulkowski et al. (2012); Proesmans et al. (2020); Scandi et al. (2022). Moreover, in many situations, the dynamical evolution of a system may show special symmetry or modularity properties that can affect its thermodynamics. In such cases, extra irreversibility may arise that can limit work extraction even when allowing generic thermodynamic processes under local detailed balance Kolchinsky and Wolpert (2021a, b).
@@
Here we will assume for simplicity no limitations associated to the relaxation time-scales, so that quasi-static processes are feasible, and no extra symmetries. We moreover focus on feedback control processes mimicking the optimal ones, i.e., consisting in the implementation of a sudden quench on the system Hamiltonian that depends on the measurement result, followed by quasi-static isothermal evolution that takes the Hamiltonian back to its original form. All the limitations considered here hence regards the set of available quenches that can be performed to extract work from the initial (nonequilibrium) state of the system by transforming it into an equilibrium state. Also, in many real-world scenarios, the experimentalist has little (if any) direct control over rate matrices, beyond that arising via local detailed balance and their ability to change a system’s energy spectrum. Accordingly (and in contrast to Kolchinsky and Wolpert (2021a, b)), we assume no control of rate matrices whatsoever beyond that which arises automatically due to local detailed balance.
The remainder of this article is organized as follows. In Sec. II we outline the framework of thermodynamic quenches in the ideal case and under control limitations. In Sec. III, we describe the interplay of measurement, action, and work extraction under constraints in several scenarios. In Sec. IV we describe our main results about optimal construction of an optimal action repertoire. In Sec. V we consider the cases of perfect observation and of a random initial distribution. In Sec. VI we provide a concluding summary and discussion.
@@ II Framework
Consider a non-equilibrium system of interest from which work is to be extracted. It is described by a set of $d_{S}$ variables $\mathbf{x}:=\{x_{i}\}_{i=1}^{d_{S}}$, with $x_{i}\in\mathbb{X}$ (each of which can be either discrete or continuous). We introduce an associated Hamiltonian function describing the energy of the system over this set of variables $H(\mathbf{x})$ which can be externally controlled by an external agent subjected to computational limitations that will be specified in detail later. The system is assumed to be initially in a non-equilibrium state described by distribution $\rho_{0}(\mathbf{x})$.
A feedback controller can extract work from the nonequilibrium state of the system by first measuring (all or some of) the system variables, and then performing a control operation over the system Hamiltonian conditioned on the obtained measurement result. As a consequence of the measurement, the controller obtains a set of results $\mathbf{m}:=\{m_{i}\}_{i=1}^{d_{M}}$ with $m_{i}\in\mathbb{M}$ (which can also be either discrete or continuous, and which can be univocally related to the system variables $x_{i}$ or not). The measurement results are affected by external noise as described by the (joint) conditional probability $P(\mathbf{m}|\mathbf{x})$ for obtaining outcome $\mathbf{m}$ given that the state of the system was $\mathbf{x}$. The (marginal) probability of obtaining the results $\mathbf{m}$ is therefore a function of the initial state $\rho_{0}(\mathbf{x})$ and the above conditional probabilities as given by
@@
as corresponds to Bayes’ rule for updating probabilities when using $\rho_{0}(\mathbf{x})$ as the prior distribution. In the case of a perfect measurement of the system variables (i.e., the measurement result $m_{i}$ correspond to the system state variable $x_{i}$ for all $i$) the above expressions reduce to $P(\mathbf{m}|\mathbf{x})=\prod_{i=1}^{d}\delta_{m_{i},x_{i}}$, where $\delta_{m,x}$ denotes the Kronecker delta, and hence $p(\mathbf{m})=\rho_{0}(\mathbf{m})$, implying that the state of the system becomes perfectly known, i.e., $\rho_{X|\mathbf{m}}(\mathbf{x})=\prod_{i=1}^{d}\delta_{m_{i},x_{i}}$. See Sec. V.1 for how our main results apply in this setting.
II.1 Ideal optimal work extraction
The maximum amount of work that can be extracted from the system, assuming knowledge of $\rho_{X|\mathbf{m}}(\mathbf{x})$, and access to an equilibrium reservoir at temperature $T$, is the difference in nonequilibrium free energies between such a state and the corresponding equilibrium state $\rho_{H}^{\rm eq}(\mathbf{x})=e^{-\beta H(\mathbf{x})}/Z$ with $Z=\sum_{\mathbf{x}}e^{-\beta H(\mathbf{x})}$ being the partition function, and $\beta=1/(k_{B}T)$, where $T$ is the temperature of the (single) thermal reservoir connected to the system Parrondo et al. (2015). That is:
$W_{\rm ext}(\mathbf{m})\leq\mathcal{F}_{H}(\rho_{X|\mathbf{m}})-F_{H}^{\rm eq},$ (3)
@@ II.1 Ideal optimal work extraction
The feedback control protocol achieving optimal work extraction (equality in Eq. (3)) can be constructed in a generic way by first instantaneously quenching the Hamiltonian to
@@
and then quasi-statically modifying from $H_{\mathbf{m}}(\mathbf{x})$ back to the original Hamiltonian $H(\mathbf{x})$, while in contact with the external bath Parrondo et al. (2015). Recall that the rationale behind this procedure follows from the fact that the initial Hamiltonian quench instantaneously transforms the state $\rho_{X|\mathbf{m}}(\mathbf{x})$ into being an equilibrium state with respect to the quenched Hamiltonian $H_{\mathbf{m}}(\mathbf{x})$, after which the quasi-static relaxation step just becomes a thermodynamically reversible transformation where the state of the system is, at every time, in equilibrium with the bath. By following this procedure, all of the available free energy is extracted. We write that available free energy as:
$W_{\rm ext}(\mathbf{m})=-\langle H_{\mathbf{m}}\rangle_{\rho_{X|\mathbf{m}}}+\langle H\rangle_{\rho_{X|\mathbf{m}}}$
@@
Notice that $\mathcal{F}_{H}(\rho_{X|\mathbf{m}})$ is a stochastic quantity depending on the specific outcomes obtained in the measurement, and hence so is $W_{\rm ext}(\mathbf{m})$. The associated average over measurement results is:
$\langle W_{\rm ext}\rangle$ $=\sum_{\mathbf{m}}p(\mathbf{m})\mathcal{F}_{H}(\rho_{X|\mathbf{m}})-F_{H}^{\rm eq}$
@@
$=k_{B}TI(\mathbf{m};\mathbf{x})+\mathcal{F}_{H}(\rho_{0})-F_{H}^{\rm eq},$ (7)
where we identified $I(\mathbf{m};\mathbf{x})=S(\rho_{0})-\sum_{\mathbf{m}}p(\mathbf{m})S(\rho_{X|\mathbf{m}})$ as the mutual information between the measurement outcomes and the system variables. In the equation above we used that the unconditional state of the system after measurement is just the initial state, $\sum_{\mathbf{m}}p(\mathbf{m})\rho_{X|\mathbf{m}}=\rho_{0}$, which can be directly verified from the expression of $\rho_{\mathbf{m}}$ in Eq. (2) (a property of efficient non-invasive measurements). Equation (II.1) tells us that the amount of work extractable from the environment is limited by two factors (i) the amount of information that can be gathered from measuring the system of interest as given by the mutual information $I(\mathbf{m};\mathbf{x})$ and (ii) the amount of free energy contained the initial non-equilibrium state of the environment $\rho_{0}(\mathbf{x})$ as compared to its equilibrium free energy.
It is worth mentioning at this point that closing the cycle after the work extraction procedure requires a final step consisting in the erasure of the information gathered during the measurement of the system and employed for the implementation of the feedback control protocol Parrondo et al. (2015). Following generalized Landauer’s principle, this erasure process has an unavoidable work cost that amounts, at least, to the mutual information $k_{B}TI(\mathbf{m};\mathbf{x})$ in the reversible limit Sagawa and Ueda (2009). That means that the net effect of the work extraction procedure, once erasure costs of the memory have been included, is to transform the free energy difference from the initial nonequilibrium state $\rho_{0}(\mathbf{x})$ to the equilibrium one $\rho_{\rm eq}(\mathbf{x})$ into useful work.

### v2

@@ Abstract
Many physical, biological, and even social systems are faced with the problem of how to efficiently harvest free energy from an environment that can have many possible states, yet only have a limited number of harvesting protocols to choose among. We investigate this scenario by extending earlier work on using feedback control to extract work from nonequilibirum systems. Specifically, in contrast to that previous work on the thermodynamics of feedback control, we analyze the combined and separate effects of noisy measurements, memory limitations, and limitations on the number of possible work extraction protocols. Our analysis provides a general recipe to construct repertoires of allowed harvesting protocols that minimize the expected thermodynamic losses during free energy harvesting, i.e., that minimize expected entropy production. In particular, our results highlight that the benefits of feedback control over uninformed (random) actions extend beyond just the associated information gain, often by many orders of magnitude. Our results also uncover the effects of limitations on the number of possible harvesting protocols when there is uncertainty about the distribution over states of the environment.
@@ I Introduction
A central concern of essentially all living organisms is that they must interact intelligently with their nonequilibrium environment in order to extract enough free energy from it to survive. This in turn requires them to acquire information concerning the state of their environment via measurements, to guide their behavior in those interactions Tkačik and Bialek (2016). Relatedly, a long line of scientists have been fascinated by the interplay between the probabilistic nature of the second law and the possibility to use information to improve thermodynamic operation Leff and Rex (2002) since Maxwell’s demon thought experiment in the 19th century and the first information engine using feedback control operations proposed by Leo Szilard Szilard (1929).
@@
The development of stochastic thermodynamics in the last decades has provided the possibility of a precise formulation and analysis of feedback control scenarios Parrondo et al. (2015) where the costs and benefits of using information can be assessed in a variety of situations, ranging from discrete measurements Touchette and Lloyd (2000); Sagawa and Ueda (2009, 2010); Horowitz and Vaikuntanathan (2010); Abreu and Seifert (2012), to continuous monitoring Sagawa and Ueda (2012); Deffner and Jarzynski (2013); Horowitz et al. (2013); Potts and Samuelsson (2018); Ribezzi-Crivellari and Ritort (2019); Debiossac et al. (2020), and from classical to quantum systems Sagawa and Ueda (2008); Funo et al. (2013); Francica et al. (2017); Manzano et al. (2018); Annby-Andersson et al. (2022). Experiments in a broad range of platforms from colloidal particles Toyabe et al. (2010) and nanoelectronic devices Koski et al. (2014a, b) to DNA pulling experiments Rico-Pasto et al. (2021) have tested and verified the role of information in work extraction.
In this paper we focus on using feedback protocols to guide the extraction of work (i.e., the harvesting of free energy) from a nonequilibrium system of interest (e.g. the environment of an organism). We consider specifically the effects arising from noisy observation and from a limited number of possible actions on the nonequilibrium system, as well as limitations on the size of the memory used in the feedback control operation. Note in particular that limitations on the number of possible actions by the agent extracting work apply in many real-world scenarios. For example, they arise in switching on and off predefined potentials Toyabe et al. (2010), suddenly raising the energy of a discrete level Koski et al. (2014a) or swapping the polarity of photons Vidrighin et al. (2016). More generally, such limitations apply to all biological organisms that are trying to harvest free energy from their environment.
Here we show that such limitations can drastically reduce the efficiency of free energy extraction performed via feedback control. As a result, determining the optimal finite set of actions the controller can choose among, given such memory limitations and noise in measurements is crucial in several applications. These include cases where the initial distribution of the system is not known and might be guessed from noisy observation. To analyze these issues, here we build on previous work demonstrating that optimal feedback control which maximizes work extraction can be achieved with reversible feedback protocols Horowitz and Parrondo (2011a, b). Such optimal control begins with a quench of the original Hamiltonian of the nonequilibrium system. The precise quench depends on the original distribution of system states and any (noise-free) measurement concerning that state. After this quench a quasi-static isothermal process is applied to the system, changing the Hamiltonian back to the original Hamiltonian, and changing the distribution to an equilibrium distribution with respect to that original Hamiltonian Hasegawa et al. (2010); Takara et al. (2010); Esposito and den Broeck (2011); Anders and Giovannetti (2013); Parrondo et al. (2015).
@@
However, as mentioned, in real-world scenarios limitations often arise due to both imperfections in the measurements and because of severe constraints in the control operations that prevent the implementation of optimal protocols Johan du Buisson and Bechhoefer (2024). For example, if the relaxation timescales of the system are not fast enough with respect to the driving velocity, the quasi-static regime is lost. In that case, optimal transport techniques are required to maximize the extractable work that usually involve numerical techniques Aurell et al. (2011). In some cases, however, the problem can be directly solved in a linear response regime if the dynamics is sufficiently slow (i.e. close to but out of the quasi-static regime) Sivak and Crooks (2012); Zulkowski et al. (2012); Proesmans et al. (2020); Scandi et al. (2022). Moreover, in many situations, the dynamical evolution of a system may show special symmetry or modularity properties that can affect its thermodynamics. In such cases, extra irreversibility may arise that can limit work extraction even when allowing generic thermodynamic processes under local detailed balance Kolchinsky and Wolpert (2021a, b).
@@
Here we will assume for simplicity no limitations associated to the relaxation time-scales, so that quasi-static processes are feasible, and no extra symmetries. We moreover focus on feedback control processes mimicking the optimal ones, i.e., consisting in the implementation of a sudden quench on the system Hamiltonian that depends on the measurement result, followed by quasi-static isothermal evolution that takes the Hamiltonian back to its original form. All the limitations considered here hence regard the set of available quenches that can be performed to extract work from the initial (nonequilibrium) state of the system by transforming it into an equilibrium state (with respect to the quenched Hamiltonian). Also, in many real-world scenarios, the experimentalist has little (if any) direct control over rate matrices, beyond that arising via local detailed balance and their ability to change a system’s energy spectrum. Accordingly (and in contrast to Kolchinsky and Wolpert (2021a, b)), we assume no control of rate matrices whatsoever beyond that which arises automatically due to local detailed balance.
Our main results involve the derivation of universal expressions for the entropy production arising due to the control limitations, which limit the amount of extractable work (free energy) from generic nonequilibrium states. The general framework is outlined in Sec. II starting from the ideal case and then introducing control limitations. That framework allows us to obtain in Sec. III precise conditions for the design of successful feedback control quenches that balance the eventual work costs with a net free energy extraction. There we also determine the quantitative benefits of employing information with respect to scenarios where the quenching Hamiltonian is fixed or taken at random from a predefined set. In Sec. IV we derive explicit expressions for the optimal form of the quenches that minimize entropy production and present our results for the construction of an optimal action repertoire. We discuss its implications in terms of the acquired information that is useful for feedback control. Our general results are particularized in various situations of interest in stochastic thermodynamics in Sec. V. Noise-free observation and observation with small errors are considered, as well as the case of a random initial distribution, all of which illustrated using worked examples. Finally, in Sec. VI we provide a concluding summary and discussion. Details about the derivations of the main results are given in Appendices A, B, C, and D, and Table 1 is a guide to the notation used in the manuscript.
@@ II Framework
Consider a non-equilibrium system of interest from which work is to be extracted. It is described by a set of $d_{X}$ variables $\mathbf{x}:=\{x_{i}\}_{i=1}^{d_{X}}$ (we use bold letters to indicate vectors), with $x_{i}\in\mathbb{X}$, where the set $\mathbb{X}$ can be either discrete or continuous. We introduce an associated Hamiltonian function describing the energy of the system over this set of variables $H(\mathbf{x})$ which can be externally controlled by an external agent subjected to computational limitations that will be specified in detail later. The system is assumed to be initially in a non-equilibrium state described by distribution $\rho_{0}(\mathbf{x})$.
A feedback controller can extract work from the nonequilibrium state of the system by first measuring (all or some of) the system variables, and then performing a control operation over the system Hamiltonian conditioned on the obtained measurement result. The measurement provides the controller with a vector of measurement results $\mathbf{m}:=\{m_{i}\}_{i=1}^{d_{M}}$ with $m_{i}\in\mathbb{M}$ (which can also be either discrete or continuous). The measurement results are affected by external noise as described by the (joint) conditional probability $p(\mathbf{m}|\mathbf{x})$ for obtaining outcome $\mathbf{m}$ given that the state of the system was $\mathbf{x}$. The (marginal) probability of obtaining the results $\mathbf{m}$ is therefore a function of the initial state $\rho_{0}(\mathbf{x})$ and the above conditional probabilities as given by
@@
as corresponds to Bayes’ rule for updating probabilities when using $\rho_{0}(\mathbf{x})$ as the prior distribution. Note that we use $\rho_{X|\mathbf{m}}(\mathbf{x})$ as a short-hand notation to denote the conditional distribution $\rho(\mathbf{x}|\mathbf{m})$ for a given measurement result $\mathbb{m}$. Analogously we will use in the following $p_{M|\mathbb{x}}(\mathbf{m})$ to denote the conditional probability $p(\mathbf{m}|\mathbf{x})$. In the case of a perfect measurement of the system variables (i.e., the measurement result $m_{i}$ corresponds exactly to the system state variable $x_{i}$ for all $i$) the above expressions reduce to $p_{M|\mathbb{x}}(\mathbf{m})=\prod_{i=1}^{d}\delta_{m_{i},x_{i}}$, where $\delta_{m,x}$ denotes the Kronecker delta, and hence $p(\mathbf{m})=\rho_{0}(\mathbf{m})$, implying that the state of the system becomes perfectly known, i.e., $\rho_{X|\mathbf{m}}(\mathbf{x})=\prod_{i=1}^{d}\delta_{m_{i},x_{i}}$. See Sec. V.1 for how our main results apply in this setting.
II.1 Ideal work extraction
The maximum amount of work that can be extracted from the system, assuming knowledge of $\rho_{X|\mathbf{m}}(\mathbf{x})$, and access to an equilibrium reservoir at temperature $T$, is the difference in nonequilibrium free energies between such a state and the corresponding equilibrium state $\rho_{H}^{\rm eq}(\mathbf{x})=e^{-\beta H(\mathbf{x})}/Z$ with $Z=\sum_{\mathbf{x}}e^{-\beta H(\mathbf{x})}$ being the partition function, and $\beta=1/(k_{B}T)$, where $T$ is the temperature of the thermal reservoir connected to the system (or bath)  Parrondo et al. (2015). That is:
$W(\mathbf{m})\leq\mathcal{F}_{H}(\rho_{X|\mathbf{m}})-F_{H}^{\rm eq},$ (3)
@@ II.1 Ideal work extraction
The feedback control protocol achieving optimal work extraction [equality in Eq. (3)] can be constructed in a generic way by first instantaneously quenching the Hamiltonian to
@@
and then quasi-statically modifying from $H_{\mathbf{m}}(\mathbf{x})$ back to the original Hamiltonian $H(\mathbf{x})$, while in contact with the external bath Parrondo et al. (2015). Recall that the rationale behind this procedure follows from the fact that the initial Hamiltonian quench instantaneously transforms the state $\rho_{X|\mathbf{m}}(\mathbf{x})$ into being an equilibrium state with respect to the quenched Hamiltonian $H_{\mathbf{m}}(\mathbf{x})$, after which the quasi-static relaxation step just becomes a thermodynamically reversible transformation where the state of the system is, at every time, in equilibrium with the bath. By following this procedure, all of the available free energy is extracted. We write the work extracted using this (ideal) feedback control protocol as:
$W_{\rm ideal}(\mathbf{m})=-\langle H_{\mathbf{m}}\rangle_{\rho_{X|\mathbf{m}}}+\langle H\rangle_{\rho_{X|\mathbf{m}}}$
@@
Notice that $\mathcal{F}_{H}(\rho_{X|\mathbf{m}})$ is a stochastic quantity depending on the specific outcomes obtained in the measurement, and hence so is $W_{\rm ideal}(\mathbf{m})$. The associated average over measurement results is:
$\langle W_{\rm ideal}\rangle$ $=\sum_{\mathbf{m}}p(\mathbf{m})\mathcal{F}_{H}(\rho_{X|\mathbf{m}})-F_{H}^{\rm eq}$
@@
$=k_{B}TI_{X;M}+\mathcal{F}_{H}(\rho_{0})-F_{H}^{\rm eq},$ (7)
where we identified $I_{X;M}:=S(\rho_{0})-\sum_{\mathbf{m}}p(\mathbf{m})S(\rho_{X|\mathbf{m}})$ as the mutual information between the system variables and measurement outcomes. In the equation above we used that the unconditional state of the system after measurement is just the initial state, $\sum_{\mathbf{m}}p(\mathbf{m})\rho_{X|\mathbf{m}}(\mathbb{x})=\rho_{0}(\mathbb{x})$, which can be directly verified from the expression of $\rho_{X|\mathbf{m}}$ in Eq. (2) (a property of non-invasive measurements). Equation (II.1) tells us that the amount of work extractable from the environment is limited by two factors (i) the amount of information that can be gathered from measuring the system of interest as given by the mutual information $I_{X;M}$ and (ii) the amount of free energy contained in the initial non-equilibrium state of the environment $\rho_{0}(\mathbf{x})$ as compared to its equilibrium free energy, which we denote in the following $\Delta\mathcal{F}_{H}:=\mathcal{F}_{H}(\rho_{0})-F_{H}^{\rm eq}$.
It is worth mentioning at this point that closing the cycle after the work extraction procedure requires a final step consisting in the erasure of the information gathered during the measurement of the system and employed for the implementation of the feedback control protocol Parrondo et al. (2015). Following generalized Landauer’s principle, this erasure process has an unavoidable work cost that amounts, at least, to the mutual information $k_{B}TI_{X;M}$ in the reversible limit Sagawa and Ueda (2009). That means that the net effect of the work extraction procedure, once erasure costs of the memory have been included, is to transform the free energy difference from the initial nonequilibrium state $\rho_{0}(\mathbf{x})$ to the equilibrium one $\rho_{H}^{\rm eq}(\mathbf{x})$ into useful work.